# PoseLess: Depth-Free Vision-to-Joint Control via Direct Image Mapping with VLM

### v1

@@ Abstract
This paper introduces PoseLess, a novel framework for robot hand control that eliminates the need for explicit pose estimation by directly mapping 2D images to joint angles using tokenized representations. Our approach leverages synthetic training data generated through randomized joint configurations, enabling zero-shot generalization to real-world scenarios and cross-morphology transfer from robotic to human hands. By tokenizing visual inputs and employing a transformer-based decoder, PoseLess achieves robust, low-latency control while addressing challenges such as depth ambiguity and data scarcity. Experimental results demonstrate competitive performance in joint angle prediction accuracy without relying on any human-labelled dataset.
@@ 1   Introduction
Recent advances in computer vision and deep learning have opened new possibilities for direct perception-to-control models that bypass intermediate pose estimation. By leveraging large-scale vision-language models (VLMs) with robust tokenization strategies, it is now feasible to map high-dimensional image inputs directly to control outputs without requiring an explicit representation of pose. This paradigm shift enables more flexible and efficient control policies that are less susceptible to the limitations of traditional approaches.
@@ 1.1   Motivation
PoseLess offers a novel alternative by leveraging a vision-language model (VLM) to tokenize visual inputs and decode them directly into joint angles. This approach not only simplifies the control pipeline but also enhances generalization across different hand morphologies, including human-to-robot transfer. By training on synthetic data with extensive domain randomization, our method eliminates the dependency on manually labeled datasets and ensures adaptability to real-world variations.
@@ 1.2   Contributions
  • We introduce a novel framework that leverages a VLM (e.g., Qwen 2.5 3B Instruct) to directly map monocular images to robot joint angles, bypassing pose estimation entirely. The VLM’s ability to ”see” and tokenize images enables robust, morphology-agnostic feature extraction, reducing error propagation inherent in two-stage pipelines.
@@ 2.2   Direct Mapping Approaches
2.3   Image Tokenization and Vision-Language Models
The use of image tokenization for robotic control is inspired by recent advancements in vision-language models (VLMs) Brohan et al. (2023); Pertsch et al. (2025). VLMs, such as those used in Palo and Johns (2024); Kim et al. (2024); Black et al. (2024), have demonstrated remarkable capabilities in understanding and reasoning about visual scenes by converting images into discrete tokens that can be processed by transformer architectures. Our approach leverages the inherent ability of VLMs Bai et al. (2025) to extract rich, contextualized visual features through tokenization, enabling a more robust and generalizable mapping from images to joint angles compared to methods relying on hand-crafted features or convolutional neural networks.
@@ 5   Discussion
The experimental results underscore the potential of the PoseLess framework as a robust alternative to traditional pose-based control. By directly mapping monocular images to joint angles through tokenized representations, our approach avoids the compounded errors typically associated with multi-stage pipelines. The observed reduction in mean squared error (MSE) up to checkpoint cp-3500 illustrates that the model effectively learns the underlying mapping between visual features and joint configurations, even when trained solely on synthetic data Levine et al. (2016).
@@
Finally, while the current implementation has achieved competitive performance, several challenges remain. Integrating real-world data into the training regime, refining the tokenization process, and extending the framework to accommodate multi-view or temporal data could provide further performance improvements. Addressing these limitations will be key to realizing the full potential of direct image-to-joint mapping in diverse and uncontrolled settings.
@@ 6   Conclusion
This paper introduced PoseLess, a novel framework for depth-free, vision-to-joint control of robotic and human hands. By leveraging the power of tokenized representations and Vision Language Models (VLMs), PoseLess successfully circumvents the need for explicit pose estimation, directly mapping monocular images to joint angles. Our experiments demonstrated that training solely on synthetic data generated through randomized joint configurations and domain randomization techniques yields competitive performance in joint angle prediction accuracy, specifically achieving a reduced mean squared error up to checkpoint cp-3500. This validates the efficacy of the poseless control paradigm in addressing challenges like depth ambiguity and data scarcity.

### v2

@@ Abstract
This paper introduces PoseLess, a novel framework for robot hand control that eliminates the need for explicit pose estimation by directly mapping 2D images to joint angles using projected representations. Our approach leverages synthetic training data generated through randomized joint configurations, enabling zero-shot generalization to real-world scenarios and cross-morphology transfer from robotic to human hands. By projecting visual inputs and employing a transformer-based decoder, PoseLess achieves robust, low-latency control while addressing challenges such as depth ambiguity and data scarcity. Experimental results demonstrate competitive performance in joint angle prediction accuracy without relying on any human-labelled dataset.
@@ 1   Introduction
Recent advances in computer vision and deep learning have opened new possibilities for direct perception-to-control models that bypass intermediate pose estimation. By leveraging large-scale vision-language models (VLMs) with robust image projection strategies, it is now feasible to map high-dimensional image inputs directly to control outputs without requiring an explicit representation of pose. This paradigm shift enables more flexible and efficient control policies that are less susceptible to the limitations of traditional approaches.
@@ 1.1   Motivation
PoseLess offers a novel alternative by leveraging a vision-language model (VLM) to project visual inputs and decode them directly into joint angles. This approach not only simplifies the control pipeline but also enhances generalization across different hand morphologies, including human-to-robot transfer. By training on synthetic data with extensive domain randomization, our method eliminates the dependency on manually labeled datasets and ensures adaptability to real-world variations.
@@ 1.2   Contributions
  • We introduce a novel framework that leverages a VLM (e.g., Qwen 2.5 3B Instruct) to directly map monocular images to robot joint angles, bypassing pose estimation entirely. The VLM’s ability to ”see” and project images enables robust, morphology-agnostic feature extraction, reducing error propagation inherent in two-stage pipelines.
@@ 2.2   Direct Mapping Approaches
2.3   Image Projection and Vision-Language Models
The use of image projection for robotic control is inspired by recent advancements in vision-language models (VLMs) Brohan et al. (2023); Pertsch et al. (2025). VLMs, such as those used in Palo and Johns (2024); Kim et al. (2024); Black et al. (2024), have demonstrated remarkable capabilities in understanding and reasoning about visual scenes by converting images into discrete tokens that can be processed by transformer architectures. Our approach leverages the inherent ability of VLMs Bai et al. (2025) to extract rich, contextualized visual features through tokenization, enabling a more robust and generalizable mapping from images to joint angles compared to methods relying on hand-crafted features or convolutional neural networks.
@@ 5   Discussion
The experimental results underscore the potential of the PoseLess framework as a robust alternative to traditional pose-based control. By directly mapping monocular images to joint angles through projected representations, our approach avoids the compounded errors typically associated with multi-stage pipelines. The observed reduction in mean squared error (MSE) up to checkpoint cp-3500 illustrates that the model effectively learns the underlying mapping between visual features and joint configurations, even when trained solely on synthetic data Levine et al. (2016).
@@
Finally, while the current implementation has achieved competitive performance, several challenges remain. Integrating real-world data into the training regime, refining the image projection process, and extending the framework to accommodate multi-view or temporal data could provide further performance improvements. Addressing these limitations will be key to realizing the full potential of direct image-to-joint mapping in diverse and uncontrolled settings.
@@ 6   Conclusion
This paper introduced PoseLess, a novel framework for depth-free, vision-to-joint control of robotic and human hands. By leveraging the power of projected representations and Vision Language Models (VLMs), PoseLess successfully circumvents the need for explicit pose estimation, directly mapping monocular images to joint angles. Our experiments demonstrated that training solely on synthetic data generated through randomized joint configurations and domain randomization techniques yields competitive performance in joint angle prediction accuracy, specifically achieving a reduced mean squared error up to checkpoint cp-3500. This validates the efficacy of the poseless control paradigm in addressing challenges like depth ambiguity and data scarcity.